OpenStudy (anonymous):

How many different combinations of Rubik's cube are there, and how can you calculate them?

Directrix (directrix):

The number of legally attainable configurations is therefore 8! * 3^8 * 12! * 2^12 --------------------- = 8! * 3^7 * 12! * 2^10 2*2*3 = 43252003274489856000 = 4.32520*10^19 (approximately) Read the discusion for this at the link: http://mathforum.org/library/drmath/view/54301.html